# Cryptanalysis on Lightweight Verifiable Homomorphic Encryption

### v2

@@ Abstract
At CCS, Chatel et al. [CKP+24] introduced two lightweight VHE schemes: Replication Encoding (REP) and Polynomial Encoding (PE). A similar approach to REP was used by Albrecht et al. [ADDG24] in Eurocrypt to develop a Verifiable Oblivious PRF scheme (vADDG). A key approach in these schemes is to embed specific secret information within HE ciphertexts to verify homomorphic evaluations.
This paper presents efficient attacks that exploit the homomorphic properties of encryption schemes. The one strategy is to retrieve the secret information in encrypted state from the input ciphertexts and then leverage it to modify the resulting ciphertext without being detected by the verification algorithm. The other is to exploit the secret embedding structure to modify the evaluation function $f$ into $f^{\prime}$ which works well on input values for verification purposes.
Our forgery attack on vADDG demonstrates that the proposed 80-bit security parameters in fact offer less than 10-bits of concrete security. Our attack on REP and PE achieves a probability 1 attack with linear time complexity when using fully homomorphic encryption.
Keywords:
Homomorphic Encryption Verifiable Computation VOPRF Cryptanalysis
@@ 1 Introduction
Verifiable Computation (VC) is a technique that guarantees the correctness of the result of a computation outsourced by a client [GGP10]. This technique allows the client to detect and prevent erroneous or malicious computations by the server. This ensures the integrity of the computations. On the other hand, Homomorphic Encryption (HE) allows computations to be performed on encrypted data without having to decrypt. HE allows the privacy of sensitive information to be maintained while complex calculations are outsourced to the server. Overall, in outsourced computing, VC ensures the integrity of the computational results, while HE guarantees the privacy of the client’s data. Verifiable Homomorphic Encryption (VHE) is the combination of these two technologies. A naive approach to VHE is the use of the SNARK or other VC techniques for homomorphic computation on encrypted data [ABPS24]. However, this approach results in a highly inefficient solution due to several complex operations required in HE, such as bootstrapping and relinearization. These operations are difficult to integrate seamlessly with existing VC techniques.
Alternatively, other lines of research [CKP+24, ADDG24] have explored the design of efficient verifiable homomorphic encryption schemes without using general VC techniques. These schemes aim to achieve verifiability at lower computational cost by concealing specific secret informations within the ciphertext and leveraging their confidentiality to ensure integrity. In this paper, we review such ‘lightweight’ verifiable homomorphic encryption schemes and demonstrate that these approaches have been unsuccessful, highlighting the need for further investigation and improvement.
@@ 1.1 Lightweight Verifiable Homomorphic Encryptions
Albrecht et al. [ADDG24] proposed a new candidate for an Oblivious Pseudorandom Function (OPRF), called the ADDG scheme, using TFHE [CGGI20]. They extended it to a Verifiable OPRF (VOPRF), called the vADDG scheme. When two parties need to jointly compute a PRF where the PRF key is privately held by one party $\mathcal{A}$ and the inputs and outputs are privately held by another party $\mathcal{B}$, homomorphic encryption provides a solution: $\mathcal{B}$ encrypts the inputs using an HE scheme and sends them to $\mathcal{A}$. Then $\mathcal{A}$ homomorphically evaluates the PRF using its own private key and sends the result to $\mathcal{B}$. Finally, $\mathcal{B}$ decrypts the received ciphertext to obtain the result.
For the verifiability extension, vADDG uses the following method: First $\mathcal{A}$ publishes $\kappa$ verification values (i.e. challenge) $\mathbf{x}^{\star}_{k}$ for $1\leq k\leq\kappa$ and their PRF evaluations $\{\mathbf{z}^{\star}_{k}\}$ with zero-knowledge proofs. If $\mathcal{B}$ wants to evaluate $\mathbf{x}_{i}$ for $1\leq i\leq\alpha$ with verification, $\mathcal{B}$ generates a vector of length $\gamma=\alpha\nu+\beta,$ consisting of randomly permuted $\nu$ copies of each $\mathbf{x}_{i}$ and $\beta$ verification values from $\{\mathbf{x}^{\star}_{k}\}_{1\leq k\leq\kappa}$. To verify the integrity of the result after outsourcing, $\mathcal{B}$ first recovers the random permutation, then checks whether the evaluation of $\mathbf{x}^{\star}_{k}$’s matches that of $\mathbf{z}^{\star}_{k}$’s. It also checks that the $\nu$ copies of each $\mathbf{z}_{i}$ all have the same value. Then $\mathcal{B}$ accepts the $\mathbf{z}_{i}$’s as honest results. The suggested parameters for achieving 80-bit security in verification are $(\alpha,\beta,\nu)=(105,10,11).$ A notable feature of this (v)ADDG scheme is that the homomorphic computation of this PRF circuit requires only one level of bootstrapping depth by removing the key-switching key.
@@
Chatel et al. [CKP+24] proposed a novel VHE scheme, called Veritas, which supports all operations in BFV and BGV. Compared to the baseline HE schemes, Veritas introduces an overhead ranging from a factor of 1 to a two-digit factor, depending on the characteristics of the circuit. The scheme’s main idea is to use random values and their precomputed results: For a given circuit, the client, acting as the verifier, holds random verification values and their precomputed results. The client then encrypts a message along with the verification values using specially designed encoders, namely Replication Encoding (REP) and Polynomial Encoding (PE). If the homomorphic computation is correct, the decrypted result will contain both the precomputed results and the desired computational result. To verify the integrity of the result, the client checks whether the computed values match the previously known values. If they do, along with additional verifications, the client accepts the result as valid.
@@
REP encodes given messages and verification values as follows: Let $n$ be a power of two. Among the slots indexed from $1$ to $n$, the challenge values $v_{i}$’s are placed in certain slots indexed by $S\subset\{1,\dots,n\}$ with $|S|=n/2$. Meanwhile, the other slots indexed by $S^{c}=\{1,\dots,n\}\setminus S$ are repeatedly filled with the message $m$. For example, if $n=4$ and $S=\{1,4\}$, the client creates a vector $(v_{1},m,m,v_{2})$ for a message $m$ and challenge values $v_{1},v_{2}$. We refer to the former as the verification slots and the latter as the computation slots. If the computation is correctly performed, the resulting vector will be $(f(v_{1}),f(m),f(m),f(v_{2}))$. The client verifies as follows: Check that all values in the verification slots match the precomputed values $f(v_{i})$’s and that the values in the computation slots are identical. If both conditions are satisfied, the client accepts the computation result as $f(m)$.
@@
PE, the other encoding method in [CKP+24], encodes a message $\mathbf{m}\in\mathbb{Z}_{t}^{N}$ as follows: the client randomly chooses $\alpha\leftarrow\mathbb{Z}_{t}^{\times}$ and a verification value $\mathbf{v}\leftarrow\mathbb{Z}_{t}^{N}$. Next, the client interpolates $\mathbf{m},\mathbf{v}$ at $Y=0$ and $Y=\alpha$, respectively. As a result, the client obtains a linear polynomial $\mathbf{m}+\left(\frac{\mathbf{v}-\mathbf{m}}{\alpha}\right)Y\in\mathbb{Z}_{t}^{N}[Y]$. Then, this polynomial is coefficient-wisely encrypted to a ciphertext polynomial $\textsf{ct}{}_{0}+\textsf{ct}{}_{1}Y\in\mathcal{R}_{q}^{2}[Y]$. The operations in $\mathcal{R}_{q}^{2}[Y]$ are performed as a polynomial in $Y$, where the operations on the coefficients are homomorphic operations. For verification, the client checks whether $\textsf{Dec}(\mathsf{F}(\alpha))=f(\mathbf{v})$, and if they match, the constant term $\textsf{Dec}(\mathsf{F}(0))$ is accepted as the desired computation result $f(\mathbf{m})$.
However, the computation cost of PE increases exponentially with the multiplication depth: squaring $\textsf{ct}{}_{0}+\textsf{ct}{}_{1}Y$ yields $\textsf{ct}{}^{\prime}_{0}+\textsf{ct}{}^{\prime}_{1}Y+\textsf{ct}{}^{\prime}_{2}Y^{2}$, and squaring it again results in $\textsf{ct}{}^{\prime\prime}_{0}+\dots+\textsf{ct}{}^{\prime\prime}_{4}Y^{4}$. This results in a significant performance degradation. To address this issue, [CKP+24] proposed the ReQuadratization (ReQ), a client-aided protocol that transforms a quartic ciphertext polynomial into a quadratic ciphertext polynomial. See Section V.D and Appendix E of [CKP+24] for more details.
@@ 1.2 Attack Description on Lightweight VHE schemes
For the attacks on these lightweight VHE schemes, the server can forge ciphertexts without knowing the secret information, but only with the secret information in the encrypted state by utilizing homomorphic computation.
@@
For the vADDG scheme, the essential step of the forgery attack is to homomorphically recover the positions in the encrypted state where identical values appear. However, since the ADDG scheme utilizes the TFHE scheme which supports only a single level of bootstrapping depth, it is difficult to carry out the forgery attack using the published verification values. To address this, we propose a method for extracting positions by evaluating a characteristic function. Specifically, the adversary can construct a characteristic function that takes each TFHE ciphertext string as input and outputs an encryption of 00 or $1$ and then adds that output to the ciphertext string itself. The forgery attack succeeds if this characteristic function outputs 00 at the verification value and outputs $1$ at one of the computation values.
Using this approach, we can construct a circuit that achieves forgery with probability approximately $2^{-\beta}=2^{-10}$ under the given 80-bit security parameter $(\alpha,\beta,\gamma)=(105,10,11)$. This probability does not change significantly with a decrease in $\alpha$ or an increase in $\nu$, and it satisfies $\lambda$-bit security only when $\beta>\lambda$.
@@
The adversary can evaluate a cheating circuit that extracts the position of a common value among the $n$ slots. For example, when $n=4$, a circuit
@@
enables the adversary to obtain the information of the verification slot in the encrypted state. We designed a simple circuit that extracts this information from the fresh ciphertext using homomorphic comparison over $\mathbb{Z}_{t}$. After recovering the verification slots in the encrypted state, the adversary can forge the computational result. For example, the following vector
@@
has the verification values $f(v_{i})$’s in the verification slots and the malicious result $g(m)$ in the computation slots, where $g$ is any other circuit different from $f$, possibly malicious.
@@
The above attack circuit utilizes homomorphic comparisons to identify the positions of repeated values. To block this attack, one could disallow comparisons by not providing the rotation key with index 1 to the server; however, this measure also prevents bootstrapping, which is not an acceptable solution in many cases. Instead, comparisons can be prevented by employing multiple secret keys. In this approach, while still granting the server sufficient computational power including bootstrapping, it prevents the adversary from executing the previous attack. We denote this variant of REP as $\textsf{REP}{}^{\sf msk}$.
However, $\textsf{REP}{}^{\sf msk}$ can also be attacked in the same manner as the attack on vADDG. By evaluating a pseudo-random characteristic function $\chi_{A}$ on each slot for a randomly chosen $A$, if a value in a slot belongs to the set $A\subset\mathbb{Z}_{t}$, then after evaluation that slot will contain 1; otherwise, it will include 0. Therefore, if $A$ contains only the message and not the verification value, this forgery attack will succeed. We present a pseudo-random characteristic function that can be implemented in BFV with a cost of $O(\log t)$. When the adversary implements this characteristic function, the attack success probability becomes greater than $(e(1+n/2))^{-1}$, which is not negligible. In contrast to the cryptanalysis on vADDG, where security could be ensured by adjusting parameters, in this case, an attack success probability of $O(n^{-1})$ always occurs regardless of the parameter choice. However, this patch transformed the deterministic attack into a probabilistic one, which allows it to operate under scenarios on a covert adversary [AL07].
@@
We now present attacks on PE assuming access to the ReQ protocol, which enables the server to perform a circuit with large multiplicative depth. In the previous attack, the secret information required for encoding, the verification slots $S$, was recovered in an encrypted state and exploited in a forgery attack. In PE, the secret information needed for encoding is $\alpha$; however, obtaining a ciphertext $\textsf{ct}{}\in\mathcal{R}^{2}_{q}$ that contains information about $\alpha$ by homomorphic computation is difficult, since even if the adversary recovers the decryption key sk, $\alpha$ is used solely to interpolate random values, which guarantees its zero-knowledge property.  [^1] However, one might regard $\textsf{Enc}(0,\dots,0)+\textsf{Enc}(1,\dots,1)Y\in\mathcal{R}^{2}_{q}[Y]$ as an encryption of $\alpha$. See Discussion in Section 7..
Nevertheless, unlike the previous case, there is another way to attack it: There exists a vulnerability in the encoding structure. Since the encoding method of PE is an interpolation at $Y=0$ and $Y=\alpha$. by leveraging the algebraic properties of $\alpha\in\mathbb{Z}_{t}$, it is possible to interpolate the desired computed ciphertexts at $Y=0$ and $Y=\alpha$ without obtaining any information about $\alpha$ or its encryption.

### v3

@@ Abstract
Chatel et al. [CKP+24] (CCS’24) introduced two VHE schemes: Replication Encoding (REP) and Polynomial Encoding (PE). A similar approach to REP was used by Albrecht et al. [ADDG24] (EUROCRYPT’24) to develop a Verifiable Oblivious PRF scheme (vADDG). A key approach in these schemes is to embed specific secret information within ciphertexts and use them to verify homomorphic evaluations.
This paper presents efficient forgery attacks against the verifiability guarantees of these VHE schemes. We introduce two attack strategies. The first targets REP and vADDG, extracting secret information in encrypted form from input ciphertexts and leveraging it to forge output ciphertexts without being detected by the verification algorithm. The second targets PE, exploiting its secret embedding structure to forge output ciphertexts that remain valid on input values for verification, yet violate the verifiability property.
Our forgery attack on vADDG demonstrates that the proposed 80-bit security parameters provide at most 10 bits of concrete security. Our attack on REP and PE achieves a probability 1 attack with linear time complexity when using fully homomorphic encryption.
@@ 1 Introduction
Verifiable Computation (VC) is a technique that guarantees the correctness of the result of a computation outsourced by a client [GGP10]. This technique allows the client to detect and prevent erroneous or malicious computations by the server. This ensures the integrity of the computations. On the other hand, Homomorphic Encryption (HE) allows computations to be performed on encrypted data without having to decrypt. HE allows the privacy of sensitive information to be maintained while complex calculations are outsourced to the server. Overall, in outsourced computing, VC ensures the integrity of the computational results, while HE guarantees the privacy of the client’s data. Verifiable Homomorphic Encryption (VHE) is the combination of these two technologies. A naive approach to VHE is the use of the SNARK or other VC techniques for homomorphic computation on encrypted data [ABPS24]. However, this approach results in a highly inefficient solution due to several non-arithmetic operations required in HE. These operations do not integrate seamlessly with existing VC techniques.
Alternatively, other lines of research [CKP+24, ADDG24] have explored the design of efficient verifiable homomorphic encryption schemes without combining VC techniques with HE schemes. These schemes aim to achieve VHE at lower computational cost by concealing specific secret information within ciphertexts and precomputed values, and using them to ensure verifiability. In this paper, we review such ‘lightweight’ verifiable homomorphic encryption schemes and demonstrate that these approaches have been insufficient, highlighting the need for further investigation and improvement.
@@ 1.1 Lightweight Verifiable Homomorphic Encryptions
An Oblivious Pseudorandom Function (OPRF) is a two-party protocol between a server $\mathcal{A}$ holding the secret PRF key $\bf k$, and a client $\mathcal{B}$ holding its private input $\bf x$. At the end of the protocol, $\mathcal{B}$ obtains $F_{\bf k}({\bf x})$, while $\mathcal{A}$ learns nothing about $x$ and $\mathcal{B}$ learns nothing about $\bf k$. HE provides a direct solution for constructing an OPRF: $\mathcal{B}$ encrypts the inputs under an HE scheme and sends the ciphertexts to $\mathcal{A}$. Then, $\mathcal{A}$ homomorphically evaluates the PRF using its private key and returns the result to $\mathcal{B}$. Finally, $\mathcal{B}$ decrypts the received ciphertext to obtain the output.
Albrecht et al. [ADDG24] leveraged this idea to propose a new candidate for an OPRF, referred to as the ADDG scheme, based on the TFHE homomorphic encryption scheme [CGGI20]. They further extended it to a Verifiable OPRF (VOPRF), which we refer to as the vADDG scheme. For the verifiability extension, vADDG adopts the following method: First, $\mathcal{A}$ publishes a set of verification values (i.e., challenges) $\{\mathbf{x}^{\star}_{k}\}_{1\leq k\leq\kappa}$ together with their PRF evaluations $\{\mathbf{z}^{\star}_{k}\}_{1\leq k\leq\kappa}$ and corresponding zero-knowledge proofs. Suppose that $\mathcal{B}$ wishes to evaluate the OPRF $F_{\mathbf{k}}$ on private inputs $\mathbf{x}_{i}$ ($1\leq i\leq\alpha$) with verification. $\mathcal{B}$ forms a tuple consisting of $\nu$ copies of each $\mathbf{x}_{i}$ and $\beta$ verification values from the pre-published set $\{\mathbf{x}^{\star}_{k}\}_{1\leq k\leq\kappa}$. This tuple, of length $\alpha\nu+\beta$, is then randomly permuted, encrypted, and finally sent to $\mathcal{A}$ and computed.
To verify the integrity of the result after outsourcing, $\mathcal{B}$ first decrypts and recovers the random permutation. Then $\mathcal{B}$ checks whether the evaluation of $\mathbf{x}^{\star}_{k}$’s matches that of $\mathbf{z}^{\star}_{k}$’s. It also checks that the copies of each $\mathbf{z}_{i}$ all have the same value. Then $\mathcal{B}$ accepts the $\mathbf{z}_{i}$’s as honest results. A notable feature of this (v)ADDG scheme is that the homomorphic computation of this PRF circuit requires only one level of bootstrapping depth by removing the key-switching key.
@@
Chatel et al. [CKP+24] proposed a novel VHE scheme, called Veritas, which supports all operations in BFV and BGV. Compared to the baseline HE schemes, Veritas introduces an overhead ranging from a factor of 1 to a two-digit factor, depending on the characteristics of the circuit. The main idea of these schemes is similar to that of vADDG scheme. For a given circuit $f$, the client holds random verification values $v_{i}$’s and their precomputed results $f(v_{i})$’s. The client then encrypts a message along with the verification values using specially designed encoders, namely Replication Encoding (REP) and Polynomial Encoding (PE). If the homomorphic computation is correct, the decrypted result will contain both the precomputed results and the desired computational result. To verify the integrity of the result, the client checks whether the computed values match the previously known values. If they do, along with additional verifications, the client accepts the result as valid.
@@
REP encodes given messages and verification values as follows: Let $n$ be a power of two. Among the slots indexed from $1$ to $n$, the verification values $v_{i}$’s are placed in certain slots indexed by $S\subset\{1,\dots,n\}$ with $|S|=n/2$. Meanwhile, the other slots indexed by $S^{c}=\{1,\dots,n\}\setminus S$ are repeatedly filled with the message $m$. For example, if $n=4$ and $S=\{1,4\}$, the client creates a vector $(v_{1},m,m,v_{2})$ for a message $m$ and challenge values $v_{1},v_{2}$. We refer to the former as the verification slots and the latter as the computation slots. If the computation is correctly performed, the resulting vector will be $(f(v_{1}),f(m),f(m),f(v_{2}))$. The client verifies as follows: Check that all values in the verification slots match the precomputed values $f(v_{i})$’s and that the values in the computation slots are identical. If both conditions are satisfied, the client accepts the computation result as $f(m)$.
@@
PE, the other encoding method in [CKP+24], encodes a message $\mathbf{m}\in\mathbb{Z}_{t}^{N}$ as follows: the client randomly chooses $\alpha\leftarrow\mathbb{Z}_{t}^{\times}$ and a verification value $\mathbf{v}\leftarrow\mathbb{Z}_{t}^{N}$. Next, the client interpolates $\mathbf{m},\mathbf{v}$ at $Y=0$ and $Y=\alpha$, respectively. As a result, the client obtains a linear polynomial $\mathbf{m}+\left(\frac{\mathbf{v}-\mathbf{m}}{\alpha}\right)Y\in\mathbb{Z}_{t}^{N}[Y]$. Then, this polynomial is coefficientwise encrypted to a ciphertext polynomial $\textsf{ct}{}_{0}+\textsf{ct}{}_{1}Y\in\mathcal{R}_{q}^{2}[Y]$. Here, $\mathcal{R}_{q}^{2}$ is the BFV ciphertext space and $\mathcal{R}_{q}^{2}[Y]$ is the set of polynomials in the variable $Y$ with coefficients in $\mathcal{R}_{q}^{2}$, and the operations on the coefficients correspond to homomorphic operations. For verification, the client checks whether $\textsf{Dec}(\mathsf{F}(\alpha))=f(\mathbf{v})$, and if they match, the constant term $\textsf{Dec}(\mathsf{F}(0))$ is accepted as the desired computation result $f(\mathbf{m})$.
However, the computation cost of PE increases exponentially with the multiplication depth: squaring $\textsf{ct}{}_{0}+\textsf{ct}{}_{1}Y$ yields $\textsf{ct}{}^{\prime}_{0}+\textsf{ct}{}^{\prime}_{1}Y+\textsf{ct}{}^{\prime}_{2}Y^{2}$, and squaring it again results in $\textsf{ct}{}^{\prime\prime}_{0}+\dots+\textsf{ct}{}^{\prime\prime}_{4}Y^{4}$. This results in an exponential performance degradation. To address this issue, [CKP+24] proposed the ReQuadratization (ReQ), a client-aided protocol that transforms a quartic ciphertext polynomial into a quadratic ciphertext polynomial. See Section V.D and Appendix E of [CKP+24] for more details.
@@ 1.2 Attack Description on Lightweight VHE Schemes
A fundamental requirement of a VHE scheme is that the server must homomorphically evaluate the designated function $f$ on the client’s ciphertexts. If the server can instead produce a valid ciphertext corresponding to another function $g\neq f$ that still passes the verification, the verifiability guarantee of the scheme is broken. In our forgery attacks on lightweight VHE schemes, the server does not need to know the embedded secret information in plaintext. Rather, these attacks leverage the capabilities of the HE schemes to obliviously recover the secret information in its encrypted form, and then use it to forge valid ciphertexts.
@@
For the vADDG scheme, the essential step of the forgery attack is to homomorphically recover the positions in the encrypted state where identical values appear. However, since the ADDG scheme utilizes the TFHE scheme which supports only a single level of bootstrapping depth, it is difficult to carry out the forgery attack using the published verification values. To address this, we propose a method for extracting positions by evaluating a characteristic function. Specifically, the adversary can construct a characteristic function that takes each TFHE ciphertext string as input and outputs an encryption of 0 or $1$ and then adds that output to the ciphertext string itself. The forgery attack succeeds if this characteristic function outputs 0 at the verification value and outputs $1$ at one of the computation values. Using this approach, we obtain a circuit that achieves a forgery probability of roughly $2^{-10}$, showing that the concrete security of the 80-bit parameter set in [ADDG24] is significantly weaker in practice.
@@
The adversary can evaluate a cheating circuit to extract the positions of common values among the $n$ slots. For instance, when $n=4$, the circuit
@@
enables the adversary to identify the verification slots in the encrypted state. We design a simple circuit that extracts this information from a fresh ciphertext using homomorphic comparison over $\mathbb{Z}_{t}$. After recovering the verification slots in the encrypted state, the adversary can forge the computational result. For example, the following vector
@@
places the verification values $f(v_{i})$ in the verification slots and the malicious values $g(m)$ in the computation slots, where $g$ is any circuit different from $f$, possibly malicious.
@@
The above attack circuit utilizes homomorphic rotations to identify the positions of repeated values across different slots. A straightforward countermeasure is to prevent such rotations by withholding the rotation key of index 1 from the server; however, this not only disables bootstrapping—often unacceptable in practice—but also depends on the heuristic assumption that the rotation key of index 1 cannot be derived from other materials. Instead, the attack can be mitigated by employing multiple secret keys, thereby preventing computations across different slots. This method allows bootstrapping to remain available, while simultaneously blocking the adversary from carrying out the previous attack. We denote this variant of REP as $\textsf{REP}{}^{\sf msk}$.
However, $\textsf{REP}{}^{\sf msk}$ can also be attacked in the same manner as the attack on vADDG. By evaluating a characteristic function on each slot for with a randomly chosen support $A$, if a value in a slot belongs to the set $A\subset\mathbb{Z}_{t}$, then after evaluation that slot will contain 1; otherwise, it will include 0. Therefore, if $A$ contains only the message and not the verification value, this forgery attack will succeed. We present a pseudorandom characteristic function with a cost of $O(\log t)$ homomorphic multiplication. When the adversary implements this characteristic function, the attack success probability becomes greater than $(e(1+n/2))^{-1}$, which is non-negligible. In contrast to the cryptanalysis on vADDG, where security could be ensured by adjusting parameters, in this case, an attack success probability of $O(n^{-1})$ always occurs regardless of the parameter choice.
@@
We now present attacks on PE assuming access to the ReQ protocol, which enables the server to perform a circuit with large multiplicative depth. In the previous attack, the verification slots $S$, the secret of REP, were recovered in an encrypted form and exploited in a forgery attack. In PE, $\alpha$ is the secret, and similarly, if $\alpha$ is recovered in encrypted form, a forgery attack can be mounted. However, it is difficult to obtain a ciphertext ct containing information about $\alpha$ through homomorphic computation, since $\alpha$ is only used for interpolating random values and hence even the plaintexts reveal nothing about $\alpha$.  [^1] However, one might regard $\textsf{Enc}(0,\dots,0)+\textsf{Enc}(1,\dots,1)Y\in\mathcal{R}^{2}_{q}[Y]$ as an encryption of $\alpha$. See Discussion in Section 7..
Nevertheless, unlike the previous case, there is another way to forge it: There exists a vulnerability in the encoding structure. Since the encoding method of PE is an interpolation at $Y=0$ and $Y=\alpha$, by leveraging the algebraic properties of $\alpha\in\mathbb{Z}_{t}^{\times}$, it is possible to interpolate the desired computed ciphertexts at $Y=0$ and $Y=\alpha$ without obtaining any information about $\alpha$ or its encryption.